72 is subtracted from 3 times a number. The result is 3 times the difference. What is the number? I'm getting Number is 36 is that correct?
Found 2 solutions by stanbon, rfer:
Answer by stanbon(75887)

72 is subtracted from 3 times a number. The result is 3 times the difference. What is the number? I'm getting Number is 36 is that correct?
---------------------
Let the number be "x":
3x-72 = 3(difference)
------------------
Question: Difference of what?
===============================
Cheers,
Stan H.

Answer by rfer(16322)

3x-72=(3x-72)3
3x-72=9x-216
-6x=-144
x=24
I get 24, I hope I set it up correctly.